OpenStudy (anonymous):

Please help! Slope intercept form 2x -3y = 12 What would this be in slope intercept form?

OpenStudy (godlovesme):

You're basically solving for y

OpenStudy (godlovesme):

@lizajune can u do that??

OpenStudy (anonymous):

no

OpenStudy (godlovesme):

Okay first step subtract 2x from both sides

OpenStudy (anonymous):

I got that far

OpenStudy (godlovesme):

2x-3y=12 2x-2x -3y = 12-2x -3y = 12-2x Do u know the last step??

OpenStudy (anonymous):

I ended up with y=15-2x Is that right?

OpenStudy (godlovesme):

Not quite u added 3 to 15 right?

OpenStudy (godlovesme):

I mean to 12?

OpenStudy (anonymous):

Yeah

OpenStudy (godlovesme):

-3 and y are multiplied so u have to do the opposite which is?

OpenStudy (anonymous):

Ok let me try again and see what I get

OpenStudy (godlovesme):

Kk (:

OpenStudy (anonymous):

y=4-2x?

OpenStudy (godlovesme):

Very close :D But u have to divide 2 by 3 as well (u can leave it as 2/3)

OpenStudy (anonymous):

So what would the slope of the line be?

OpenStudy (godlovesme):

Y=-2/3x+4 ^^ is written in the slope-intercept form which is y=mx+b M is slope B is y-intercept

OpenStudy (anonymous):

Thanks a million!

OpenStudy (godlovesme):

Slope is -2/3 Y-intercept is 4

OpenStudy (godlovesme):

You're very Welcome!(: